OpenStudy (anonymous):

Name the variables, amplitude, period, phase shift, right end point, x and y axis orientation, and the vertical shift for f(x)=3cos(pi x/2+pi/2)-2.

OpenStudy (anonymous):

To clarify, is this your equation? \[3 \cos (\frac{ \pi \times x }{ 2 } + \frac{ \pi }{ 2 }) -2\]

OpenStudy (anonymous):